In an electroplating process, copper (ionic charge +2e, atomic weight 63.6 g/mol) is deposited using a current of 10.0 A. What mass of copper is deposited in 10.0 minutes? Avogadro's number is 6.022
Answer #1

Here Charge passes, Q = It = 10*10*60

= 6000 C

1 atom require charge to Deposited = 2*1.6*10^-19

Therefore

No. of Atoms Deposited = (6000)/(2*1.6*10^-19)

= 1.875*10^22 atoms

Therefore

No. of Moles Deposited = (1.875*10^22)/(6.023*10^23)

= 0.03113 moles

1 mole of Copper has mass = 63.5 g

Therefore

Mass of Copper Deposited = 63.5*0.03113

= 1.98 g

So E option is Correct.